### v1

# MultiVox: Benchmarking Voice Assistants for Multimodal Interactions

Abstract
The rapid progress of Large Language Models (LLMs) has empowered omni models to act as voice assistants capable of understanding spoken dialogues. These models can process multimodal inputs beyond text, such as speech and visual data, enabling more context-aware interactions. However, current benchmarks fall short in comprehensively evaluating how well these models generate context-aware responses, particularly when it comes to implicitly understanding fine-grained speech characteristics, such as pitch, emotion, timbre, and volume or the environmental acoustic context such as background sounds. Additionally, they inadequately assess the ability of models to align paralinguistic cues with complementary visual signals to inform their responses. To address these gaps, we introduce MultiVox, the first omni voice assistant benchmark designed to evaluate the ability of voice assistants to integrate spoken and visual cues including paralinguistic speech features for truly multimodal understanding. Specifically, MultiVox includes 1000 human-annotated and recorded speech dialogues that encompass diverse paralinguistic features and a range of visual cues such as images and videos. Our evaluation on 9 state-of-the-art models reveals that, although humans excel at these tasks, current models consistently struggle to produce contextually grounded responses. Our benchmark will be open-sourced and we have anonymously put examples here.  [^1] https://sites.google.com/view/multivox/
MultiVox: Benchmarking Voice Assistants for Multimodal Interactions
Ramaneswaran Selvakumar, Ashish Seth, Nishit Anand, Utkarsh Tyagi, Sonal Kumar, Sreyan Ghosh, Dinesh Manocha University of Maryland, College Park {ramans, aseth125, dmanocha}@umd.edu
1 Introduction
With recent advancements in Multimodal Large Language Models (MLLMs) (Xu et al., 2025; Microsoft et al., 2025), there is a growing interest in developing models that can understand and generate information across multiple modalities, such as images, video, and audio-simultaneously. This evolution is paving the way for the development of Omni Language Models (OLMs), which are crucial for building efficient and versatile Artificial General Intelligence (AGI) (Bubeck et al., 2023; Morris et al., 2024). While OLMs provide a wide range of applications (Xu et al., 2025), one of their primary use cases is developing omni-modal voice assistants (Huang et al., 2024). Unlike traditional speech dialogue models such as Alexa or Siri that rely solely on speech instruction, OLM-powered voice assistants (OVAs), such as GPT-4o (OpenAI, 2024) and Qwen2.5 Omni (Xu et al., 2025), can understand speech dialogues and reason over multimodal inputs, including images and videos.
Figure 1: Comparison of existing benchmark with MultiVox. Existing benchmark for omni-modal voice assistant evaluation are derived from vision-centric text VQA benchmarks. In MultiVox, models need leverage not only visual cues but also non-verbal speech signals.
Advancing OLMs as voice assistants poses challenges not only in model development but also in constructing effective evaluation benchmarks. While existing OLM benchmarks like OmniBench (Li et al., 2024) incorporate multimodal inputs such as images and video, they lack spoken dialogues—an essential modality for assessing the conversational and auditory capabilities required of voice assistants. On the other hand, current voice assistant benchmarks such as VoXDialogue (Cheng et al., 2025), SD-Eval (Ao et al., 2025), and S2S-Arena (Jiang et al., 2025) focus primarily on evaluating a model’s ability to generate contextually appropriate responses based on auditory cues like content, emotion, or speaker demographics embedded in a speech instruction. However, these benchmarks fall short of capturing the full multimodal reasoning abilities expected of OVAs, particularly in integrating visual cues alongside speech instructions.
To address this gap, we introduce MultiVox, a novel benchmark designed to evaluate an OVAs ability to incorporate multimodal cues to provide accurate and contextual responses. MultiVox includes 1000 questions consisting of human spoken questions paired with either a video or an image. Unlike existing benchmarks which primarily test visual grounding and use speech to deliver the content of a straightforward instruction, MultiVox consists of questions which require a model to combine visions skills such as object recognition, scene understanding, scene text understanding with speech skills such as acoustic scene understanding, paralanguage understanding and speaker profiling (See fig. 1). The spoken questions in MultiVox are recorded by professional voice actors to cover a diverse range of paralinguistic and emotional features that are not possible with current text-to-speech systems. A key problem in benchmarks that evaluate multi-modal reasoning capability is that models can take shortcuts by exploiting priors from other modalities, to mitigate this we introduce confounding samples in MultiVox. Specifically, each question in our benchmark has another associated question which has the same textual and visual content but their speech property is flipped such that expected answers should be different. Our key contributions are:
  1. 1.
    We present MultiVox, the first benchmark designed to evaluate omni-modal voice (OVAs) assistants using human-spoken queries paired with visual inputs. The 1000 examples require models to jointly ground visual and paralinguistic speech cues to produce accurate, context-aware responses.
  2. 2.
    We evaluate 9 omni-modal models on MultiVox and find that, while humans excel with ease, current OVAs consistently struggle, particularly with grounding speech signals, revealing a critical bottleneck in their capabilities.
  3. 3.
    We perform extensive qualitative and quantitative analysis on model’s responses and uncover key insights: Models exhibit strong visual grounding but rely heavily on textual cues for speech-related tasks; they often ignore non-verbal audio signals like tone or background sounds.
Figure 2: Illustration of various types of questions in MultiVox. We broadly define three categories of speech-understanding skills that a voice assistant needs in order to provide an accurate and contextual response. Each question in MultiVox has a speech confounder, where the textual question remains same but the speech property is flipped to counter the possibility of models exploiting unimodal priors
2 Related Work
Omni Voice Assistants The recent development of Omni Language Models (OLMs) has enabled development of omni-modal voice assistants that can simultaneously infer across both visual and speech inputs. Recent iterations of previously mentioned voice assistants now support visual inputs in form of either images like in the case of Mini-Omni2 (Xie and Wu, 2024) and MoshiVis (Royer et al., 2025) or video inputs such as Qwen2.5-Omni (Xu et al., 2025). While these models demonstrate impressive instruction-following capabilities as voice assistants, when extensively evaluated on MultiVox, we find that they often overlook crucial paralinguistic cues—such as tone, emotion, and pitch—in speech input, which are essential for generating context-aware responses.
Benchmarks For Voice Assistants While there are works such as OmniBench (Li et al., 2024) and OmniXR (Chen et al., 2024a) that evaluate OLMs, there have been few efforts to standardize the evaluation of omni voice assistants. Recent work such as Lyra (Zhong et al., 2024) repurpose existing visual question answering (VQA) benchmarks by converting the textual questions to speech. However, such approach overlooks crucial non-verbal information typically present in spoken conversations. While some progress has been made in evaluating speech VAs, many existing benchmarks still fall short. VoiceBench (Chen et al., 2024b) assesses capabilities like world knowledge and instruction following by converting textual benchmarks like MMLU (Hendrycks et al., 2021) and AlpacaEval (Dubois et al., 2024) to speech, but overlooks the non-verbal speech information. SD-Eval (Ao et al., 2025) is a pioneering work in evaluating paralinguistic features but is limited to only four categories with a narrow range of labels in each category. While VoxDialogue (Cheng et al., 2025) covers more diverse speech attributes, it relies heavily on synthetic speech generated through TTS systems that struggle to accurately convey emotional nuances and prosodic variations present in natural human speech.
3 MultiVox
We introduce MultiVox, a novel benchmark for evaluating omni-modal voice assistants on their ability to jointly interpret speech and visual inputs, and integrate them with world knowledge and reasoning to produce contextually appropriate responses. This section outlines the benchmark’s design goals, construction process, and key dataset statistics.
Figure 3: Data Curation And Annotation Pipeline Of MultiVox. The pipeline begins with task design and a taxonomy of speech traits, followed by expert annotation using multimodal sources. Professionally recorded prompts undergo quality assessment, with final review guiding benchmark construction.
3.1 Benchmark Design
Motivation Real-world communication is inherently multimodal, combining what is said with how, by whom, and in what environment. To achieve human-like understanding, an omni-modal voice assistant (OVA) must jointly interpret visual and auditory inputs—not just for content, but for paralinguistic cues such as tone, emotion, and acoustic context.
While recent progress has advanced visual understanding, speech remains a key bottleneck. Most benchmarks treat spoken input as text, overlooking critical non-verbal signals. Yet in real interactions, these cues determine how instructions are interpreted. For example, a user asking “Am I being too loud?” in a library depends on speech volume, not words alone.
To address the limitations of current benchmarks, MultiVox evaluates OVAs across three core skill domains—with a particular emphasis on speech grounding, which is an underdeveloped area in existing omni-modal benchmarks.
Speech Skills The ability to infer non-verbal attributes from speech, such as emotion, background sounds, speaker age, or tone—beyond what is conveyed in textual content.
Vision Skills The ability to recognize and interpret visual elements in images or videos, including object identity, scene understanding, reading scene text etc.
General Skills The ability to integrate visual and auditory inputs with world knowledge and common sense to make contextually appropriate decisions.
Confounder Pairs To test whether models genuinely ground non-verbal speech attributes—rather than relying on linguistic or visual shortcuts—we introduce confounder pairs. Each pair consists of two instances that share identical textual content and visual input, but differ in a targeted speech property (e.g., tone, emotion, background sound). Crucially, this difference is sufficient to invert the expected answer.
For example, a user asking “Am I being too loud?” in a library setting may expect reassurance when spoken softly, but concern or correction when spoken in a normal tone. By controlling for all other modalities, confounder pairs isolate the model’s ability to interpret non-verbal speech signals.
3.2 Benchmark Construction
Task Definition Our benchmark is organized around a three-level taxonomy of speech-related categories. The top level defines three domains: acoustic scene, speaker profile, and paralanguage. We then worked with eight expert annotators (graduate students in speech processing) to expand each domain into:
  • Categories: High-level categories of non-verbal speech attributes (e.g., emotion, speaker age, ambient environment)
  • Sub-categories: Fine-grained, measurable skills within each group which can be specified as a measurable task (e.g., detect emotion, estimate age as elderly or young adult).
For each task, experts authored a specification card detailing the speech attribute being evaluated. These cards included a clear definition of the target attribute, illustrative scenarios combining speech and vision, and structured guidelines for how annotators should construct samples and determine correct answers.
Expert Annotation Each expert was assigned a subset of tasks and followed a standardized pipeline to create benchmark samples. For each task, annotators first constructed a realistic scenario in which an OVA must correctly interpret the target speech attribute. They also created a corresponding confounder scenario—identical in text and visual input but differing in the relevant speech property—to ensure robust grounding.
To pair each question with appropriate visual content, annotators were given access to stock media libraries, prioritizing real videos. If no suitable match was found, AI-generated videos were used, though these were often insufficient for scenes involving text or fine-grained detail. In such cases, high-quality images—either retrieved or generated—were used instead. For tasks involving ambient audio (e.g., acoustic scenes), annotators selected relevant background sounds or music from curated sound libraries. In the end we collect 560 images and 440 videos.
Each finalized sample includes a textual query, accompanying visual input (image/video), and any required background audio. For paralanguage-based tasks, annotators also specified detailed voice delivery guidelines to guide later voice recording.
Reference Answers Annotators also authored reference answers and rationales for each sample, describing the expected model behavior and explaining how the relevant speech (and visual) cues should inform the response. These were later used for evaluation.
Voice Recording We employed professional voice actors to record the spoken queries, providing delivery guidelines based on the target speech property. To preserve authenticity, actors were given creative freedom in expression as long as the intended cue was conveyed. A professional audio engineer handled background sound overlays where applicable.
Quality Control Two annotators independently verified whether the intended speech attribute was clearly conveyed in each recording. Recordings without unanimous approval were revised based on feedback—either through re-recording or by adjusting the script to better support the desired delivery.
Expert Analysis In the final review stage, annotators assessed each completed sample for overall quality. They were instructed to verify that (1) the scenario was realistic, (2) the sample minimized reliance on language or vision priors, and (3) the recorded speech clearly conveyed the intended attribute.
3.3 Evaluation Criteria
The goal of MultiVox is to assess how well omni-modal voice assistants (OVAs) integrate speech and vision to produce contextually grounded responses. To enable fine-grained diagnosis, we adopt a modular evaluation framework that tests both multimodal integration and unimodal grounding. Each benchmark sample is designed to probe one or more of the following components:
Speech Grounding Evaluates whether the OVA correctly interprets paralinguistic cues such as emotion or ambient sound. Each question contains a speech hook, a non-verbal attribute that is critical for answering correctly. These tasks help isolate model sensitivity to speech signals beyond text.
Visual Grounding Evaluates whether the OVA can interpret and incorporate key visual cues. Each sample includes a visual hook—a necessary visual detail (e.g., object, background element) that the model must recognize to respond appropriately.
Contextual Appropriateness We adopt appropriateness (Chen et al., 2023) as our core evaluation metric which measures how well an OVA produces a response that aligns with the intent, context, and modality of the input.
We evaluate contextual appropriateness using sample-specific rubrics that guide judgment based on three elements: (1) a reference answer authored by the expert annotator, (2) a short rationale explaining what cues are necessary to arrive at the correct answer, and (3) task-level metadata specifying which modality is critical (e.g., speech hook, visual hook). These are provided to a GPT-4 judge, which scores model responses on a 1–5 scale, reflecting increasing levels of multimodal integration and contextual fidelity. Note that we are still experimenting with the evaluation judge and would like to switch to an open-source LLM for evaluation.
3.4 Comparison With Other Benchmark
Type Name Vis. Para. Src. Conf.
Foundation AudioBench Human
Foundation MMAU Mixed
Foundation OmniBench Human
Chat SD-Eval Human
Chat VoxDialog Synthetic
Chat S2S-Arena Mixed
Chat Lyra SVQA Synthetic
Chat Ours Human
Table 1: Comparison of MultiVox with related benchmarks.
In this section we highlight how MultiVox is different in terms of question types, modality coverage, speech source, and diagnostic power. Table 1 summarizes these differences.
Chat-based Questions Benchmarks such as AudioBench (Wang et al., 2025), MMAU (Sakshi et al., 2024), and OmniBench (Li et al., 2024) primarily test foundational tasks, with only the latter supporting full omni-modality. In contrast, MultiVox is grounded in chat-style interaction, reflecting how OVAs are deployed in real-world use. This setting demands deeper contextual understanding and flexible reasoning.
Name Size Acoustic Scene Paralanguage Speaker Profile Avg. CA
VG SG CA VG SG CA VG SG CA
Human - 90.30 82.50 4.37 87.5 92.5 4.33 94.50 95.10 4.36 4.35
Open Source Models
Mini Omni2 7.0B 78.50 20.45 1.25 75.50 28.40 1.38 83.80 9.00 1.47 1.35
VITA 1.5 1.6B 83.20 23.55 2.38 79.60 31.40 2.63 88.80 12.00 3.14 2.66
Qwen 2.5 Omni 7.0B 84.87 24.75 2.92 82.10 33.20 2.42 90.60 13.10 3.38 2.91
Baichuan Omni 7.0B 78.90 22.05 1.21 77.20 31.00 1.42 83.30 10.20 1.51 1.35
Mini CPM 8.0B 81.90 22.65 2.16 78.80 30.40 1.62 87.90 10.80 2.52 2.11
Intern Omni 8.7B 83.60 22.85 2.41 80.46 31.90 2.36 89.69 11.40 3.06 2.58
Lyra 9.0B 83.72 23.35 2.39 80.65 31.6 2.33 89.42 11.30 3.02 2.55
phi4 multimodal 5.6B 84.87 24.15 2.73 82.10 32.3 2.81 90.60 12.60 2.87 2.79
Proprietary Models
Gemini 2.0 Flash - 85.10 24.70 2.85 81.75 33.05 2.93 90.25 13.08 2.98 2.91
Table 2: Performance breakdown of human and model responses on MultiVox across key skill domains. Visual Grounding (VG) and Speech Grounding (SG) evaluates the ability to perceive specific information in the modality needed for answering the question. Contextual Appropriateness (CA) evaluates the ability to produce contextually appropriate and accurate answers given the multimodal cues
Multi-Modal Inputs Benchmarks like SD-Eval (Ao et al., 2025), VoxDialog (Cheng et al., 2025), and S2S-Arena focus on speech-only inputs, targeting paralanguage or acoustic scene understanding in isolation. Others like Lyra SVQA (Zhong et al., 2024) incorporate visual input but neglect paralinguistic cues. MultiVox is unique in requiring models to jointly interpret speech, vision, and background context, aligning with the OVA use case.
Human Speech Most existing chat benchmarks rely on synthetic speech, which current TTS systems struggle to render with accurate emotion or prosody (Wu et al., 2024; Tang et al., 2023). In MultiVox, all spoken queries are recorded by professional voice actors to preserve natural paralinguistic signals. To validate this, we conduct a user study across 100 paralanguage-focused samples, comparing human-recorded speech to CosyVoice (Du et al., 2024) and ElevenLabs TTS (ElevenLabs Inc., 2025). Ten annotators rated speech on (1) attribute match and (2) naturalness. Human speech scored 4.6/4.5, compared to 2.4/2.1 (CosyVoice) and 3.1/3.3 (ElevenLabs), supporting the decision to use professional recordings.
Confounders Many existing benchmarks are susceptible to shortcut exploitation via textual or visual priors (Kiela et al., 2021; Goyal et al., 2017). MultiVox introduces confounder pairs, where paralinguistic speech properties are inverted while keeping textual and visual input constant. This isolates the model’s ability to process non-verbal speech information and provides a more diagnostic and fine-grained evaluation framework.
4 Experiments
Evaluated Models We evaluate a wide range of OVAs. The proprietary model assessed is Gemini-2.0-flash (Developers, 2024).For open-source OVAs we use Mini-Omni2 (Xie and Wu, 2024), MiniCPM-o2.6 (Hu et al., 2024), Phi4-MM (Microsoft et al., 2025), VITA1.5 (Fu et al., 2025), Baichuan Omni (Li et al., 2025), Intern Omni (Chen et al., 2024c) and Lyra (Lee et al., 2023).
5 Main Results
Table 2 summarizes the performance of several proprietary and open-source OVAs on MultiVox. We highlight three key findings:
  • MultiVox is challenging. Although the tasks are straightforward for humans (average CA score: 4.5), the best-performing model (Gemini) only achieves 2.9—indicating that current OVAs struggle to integrate multimodal cues, particularly non-verbal speech signals, even in seemingly simple scenarios.
  • No significant gap between proprietary and open-source models. The open-source model Qwen2.5Omni matches the performance of the proprietary Gemini, though both achieve relatively low scores, highlighting that current OVAs remain limited in their ability to handle multimodal, speech-grounded tasks.
  • Speech grounding remains the bottleneck. All models show relatively strong visual grounding, but consistently struggle to interpret non-verbal speech cues such as tone, emotion, and background sounds.
To better understand the sources of these limitations, we conduct a detailed analysis of Gemini—the best-performing model on MultiVox. We examine its behavior across core skill domains, focusing on how well it grounds responses in visual and speech cues, and identifying the types of errors that arise.
Figure 4: Distribution of error types in Gemini’s responses on the MULTiVOX benchmark. The majority of errors (68.3%) stem from speech perception, indicating difficulty in interpreting non-textual auditory cues.
Figure 5: Distribution of model responses across confounder pairs. Bars indicate whether model answers were the same or different when speech cues were flipped. Ideally, answers should differ (left bar), but models often give identical responses (right bar), showing insensitivity to non-verbal speech cues.
5.1 Where do models fall short?
We analyze Gemini’s performance across three skill domains: visual grounding, speech grounding, and multimodal reasoning. This breakdown highlights where current OVAs are reliable—and where they still fall short.
5.1.1 Visual Grounding: A Strength
Table 2 shows the vision grounding scores for Gemini. We find that the model demonstrates consistently strong visual grounding capabilities. Across object detection, scene understanding, and scene text recognition, it achieves high grounding scores of 85.84, 83.92, and 87.45, respectively. These results indicate robust understanding of both fine-grained visual elements and broader scene context. Most errors arise from under-specificity in open-ended scenarios - for instance, recognizing a “person holding game controllers” without identifying attributes like “retro game controllers”. This level of ambiguity is expected in real-world scenes and the high performance even in such conditions represents strong visual grounding capabilties.
5.1.2 Speech Grounding: A Bottleneck
In contrast to visual grounding strengths, our analysis reveals substantial weaknesses in speech grounding capabilities (Table 2).
Models hear the voice, but don’t ’see’ the speaker We evaluated the model’s ability to infer demographic characteristics (age and gender) from speech cues. Results show poor performance with a speech grounding score of only 13%. Analysis of the responses show that the model relies on textual rather than acoustic acoustic content to identify speaker attributes in 80% of cases. Additionally, the model rarely commits to definitive answers (15.6% of cases), instead offering ambiguous answers (31.6%), expressing uncertainty (35.2%), or refusing to respond (17.6%).
To better understand the model’s behavior, we analyze the performance across different speech attributes. For questions requiring gender inference, the model provides neutral responses in 51% of cases, with the remaining responses showing no significant bias toward either male or female speakers. Notably, when the model does make a gender inference, it appears to be influenced primarily by visual context rather than speech characteristics. Age-related inferences reveal a stronger bias pattern, with responses overwhelmingly favoring young-to-middle-aged adult classifications regardless of the speaker’s actual age. When paired with children’s voices, the model correctly identified speakers as children in only two instances, both times due to supporting visual evidence rather than acoustic features. Similarly, elderly speakers were rarely recognized as such, with identification occurring almost exclusively when visual cues reinforced age-related characteristics. These errors have functional implications. For instance, the model suggested potentially unsafe activities to speakers with children’s voices in 20% of cases.
Models Struggle to Recognize Background Sounds In this category we test the understanding of background music, sound and ambient noise. The evaluation of acoustic scene understanding reveals significant limitations, with the model achieving only a 23.7% grounding score in this category. The model correctly identifies background sounds in just 28% of cases, with particularly poor performance in music recognition compared to general environmental sounds. Of the 28% correctly identified cases, general background sounds account for approximately 89%, while music recognition constitutes only about 11% of successful identifications. In questions with ambient noise, we observe that the model demonstrates resilience to noise when processing queries. However, it still does not have the capability to perceive noise. It appears to rely predominantly on visual cues rather than acoustic features, as evidenced in approximately 17% of cases where noisy environments were identified primarily through visual context (such as crowded airports).
Models rely on textual cues for emotion understanding Our analysis reveals that in 71% of cases, the model used textual content rather than acoustic features to infer emotion. While emotions typically have a strong correlation with text, we deliberately introduced adversarial samples (Fig 2 where textual and acoustic emotional cues conflict, leading to the model correctly identifying emotion in only 48% of instances. We observe similar performance trends in other paralanguage categories where there is an over-reliance on textual cues.
5.2 What causes these errors?
We conducted a manual error analysis of Gemini’s outputs on MultiVox to identify underlying failure patterns. Fig. 5 shows that perception errors dominate, accounting for 68.3% of all failures, primarily reflecting the model’s inability to ground to speech cues. Reasoning failures constitute 15.6% of errors, indicating that even when the model successfully perceives multimodal inputs, it struggles to effectively integrate this information to generate appropriate responses. Instruction understanding failures represent a similarly significant error category, where the model defaults to describing visual content rather than addressing the intended query. The overwhelming majority of speech perception errors indicates a clear bottleneck: improving speech perception capabilities is essential for building effective OVAs in multimodal contexts.
5.3 Do models really listen to speech cues?
To assess whether models are truly leveraging speech cues, we analyze their responses across confounder pairs, focusing on the top two performers: Gemini and Qwen2.5-Omni (Fig 5). Without considering confounders, both models appear moderately accurate—getting around 50% of responses correct (Gemini: 47%, Qwen: 51%) across these pairs. However, this aggregate accuracy can be misleading. When we examine whether models actually change their answers in response to flipped speech cues, we find that in over 45% of the confounder pairs, the model outputs are identical, i.e., they show NO FLIP in answer. Crucially, within those NO FLIP cases, most instances with one correct answer (77%) suggest that the correctness arises from chance or visual/textual bias—not from grounding in speech. This indicates that, despite non-trivial accuracy, models are largely ignoring non-verbal speech cues when answering.
6 Conclusion
We introduce MultiVox, the first benchmark designed to evaluate Omni Language Models (OLMs) as omni-modal voice assistants (OVAs) that integrate speech and vision for context-aware reasoning. Unlike prior benchmarks that rely on synthetic speech or focus only on unimodal cues, MultiVox includes 1000 professionally recorded, human-spoken questions paired with images or videos, emphasizing paralinguistic signals like tone, emotion, and background noise. A key innovation is the use of confounder pairs—speech variants with identical text and visuals—to ensure models attend to speech beyond surface cues. MultiVox enables fine-grained diagnosis across speech, vision, and general reasoning skills. Evaluation of 9 state-of-the-art models shows that, while visual grounding is robust, speech grounding remains a significant bottleneck. Our benchmark will be open-sourced to support the development of truly multimodal voice assistants.
Limitations
  • We limit ourself to questions to the english language, extending to multilingual settings is an important future direction to assess OLM generalization across languages.
  • In this work, we limit our evaluation to the content of the speech outputs and not the speech quality of the output, such as naturalness and appropriateness. Evaluating speech synthesis and conversational prosody is an important but orthogonal direction left for future benchmarks.